OpenStudy (anonymous):

Which is the equation of the given line in standard form? (Points : 1) –8x + 9y = 23 –8x + 9y = –23 –8x + 7y = 25 –9x + 8y = –23

OpenStudy (anonymous):

@math&ing001

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Standard form is Ax + By = C

OpenStudy (anonymous):

ok what is a,b, c

OpenStudy (anonymous):

Numbers..we have to find out what it is in Slope-Intercept form first, then we can re-arrange it to Standard Form. First, we find the slope. We are given two points, (-4, -1), and (0.5, 3) Plug it in the slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{3+1}{0.5+4}\) Can you add those?

OpenStudy (anonymous):

4/ 4.5

OpenStudy (anonymous):

Yep, now divide.

OpenStudy (anonymous):

1.125?

OpenStudy (anonymous):

No..well we can change 4/4.5 to 8/9.

OpenStudy (anonymous):

So the slope is 8/9.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Ugh, forget it. I'll just graph all of them. https://www.desmos.com/calculator/0r0fvdtcnh

OpenStudy (anonymous):

So what's your answer?

OpenStudy (anonymous):

the red one?

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

You got it!

OpenStudy (anonymous):

yay! can u help with couple more?

OpenStudy (anonymous):

Which is the equation of the given line? (Points : 1) x = –4 y = –4 y = –4x x = 4

OpenStudy (anonymous):

/Users/Leezo/Desktop/mediaasset_1312644_1.jpg

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Sure, but make a new post.